Steven E Jens

20 Lee Street Cambridge, MA 02139 617-864-3311 sjens@bu.edu

Objective

An entry-level position in investment management.

Education

BU School of Management Boston, Massachusetts Candidate for Master’s of Science in Investment Management (2004). GMAT: 780 Current classes include Accounting, Finance, and Quantitative Foundations. The Massachusetts Institute of Technology Cambridge, Massachusetts BS in Mathematics with Computer Science with minors in Economics and Mechanical Engineering (1998). Classes included Econometrics, Stochastic Processes, Advanced Macroeconomics, Game Theory, Algorithms, Computation Structures, Advanced Algorithms, and Cryptography.

Skills

Financial Skills/Background Enrolled in the AIMR’s CFA program. Have managed my own accounts since 1998, with better-than-market results. Computer Skills Unix (Solaris and Linux), HTML, emacs, basic makefiles, X12 EDI, and various Internet protocols. Languages include Java, Perl, Postscript, C, C++, Emacs Lisp, and M/MUMPS.

Professional Experience

Taxi Driver 2001 - 2003 (part-time) During this time period, I also spent time studying for the CFA, running for city council, and assisting a candidate for Lieutenant Governor. USPowerSolutions LLC, Java Developer 1998-2001 Wrote software to assist energy suppliers in communication with utilities. I used such technologies as EJBs (on WebLogic) and JSP to some extent, though mostly straight Java. InterSystems Corporation, Object group 1997-1998 (part-time during the school year) Wrote and tested a feature to convert M/MUMPS (a programming language produced by InterSystems) objects into Java. Athena Consulting, MIT Information Services 1996-1997 Answered questions users had about MIT’s main computer network and the software thereon. This consisted primarily of communicating what I knew to people who knew substantially less, not only about the particular system, but about computers in general.